# Safe RLHF-V: Safe Reinforcement Learning from Human Feedback in Multimodal Large Language Models

### v1

@@ Abstract
Multimodal large language models (MLLMs) are critical for developing general-purpose AI assistants, yet they face growing safety risks. How can we ensure that MLLMs are safely aligned to prevent undesired behaviors such as discrimination, misinformation, or violations of ethical standards? In a further step, we need to explore how to fine-tune MLLMs to enhance reasoning performance while ensuring they satisfy safety constraints. Fundamentally, this can be formulated as a min-max optimization problem. In this study, we propose Safe RLHF-V, the first multimodal safety alignment framework that jointly optimizes helpfulness and safety using separate multimodal reward and cost models within a Lagrangian-based constrained optimization framework. Given that there is a lack of preference datasets that separate helpfulness and safety in multimodal scenarios, we introduce BeaverTails-V, the first open-source dataset with dual preference annotations for helpfulness and safety, along with multi-level safety labels (minor, moderate, severe). Additionally, we design a Multi-level Guardrail System to proactively defend against unsafe queries and adversarial attacks. By applying the Beaver-Guard-V moderation for 5 rounds of filtering and re-generation on the precursor model, the overall safety of the upstream model is significantly improved by an average of 40.9%. Experimental results demonstrate that fine-tuning different MLLMs with Safe RLHF can effectively enhance model helpfulness while ensuring improved safety. Specifically, Safe RLHF-V improves model safety by 34.2% and helpfulness by 34.3%. All of datasets, models, and code can be found at https://github.com/SafeRLHF-V to support the safety development of MLLMs and reduce potential societal risks.
@@ 1 Introduction
Multimodal large language models (MLLMs) are pivotal for developing general-purpose AI assistants [20, 40]. By using visual instruction tuning on foundation models [15], MLLMs can effectively handle complex multimodal tasks. However, as MLLMs continue to advance, emerging studies reveal a critical challenge: How can we ensure that MLLMs are safely aligned to prevent undesired behaviors such as discrimination, misinformation, or violations of ethical standards? Additionally, recent research highlights that images can implicitly induce MLLMs to generate harmful content, exposing vulnerabilities that are less pronounced in purely language-based models. What is worse, fine-tuning MLLMs can lead to the forgetting of safety alignment previously learned by the backbone LLMs [41]. Consequently, improving the safety of MLLMs to align with human intention and values is a major concern at present [11].
@@
Reinforcement Learning from Human Feedback (RLHF) has proven effective in aligning LLMs with human preferences [21, 27], enhancing capabilities in instruction following, reasoning, etc. However, balancing helpfulness and safety remains challenging. Specifically, during the preference annotation stage, it is challenging to effectively account for two conflicting metrics within a single partial order. Helpfulness and safety are inherently somewhat contradictory, for instance, simply refusing to answer may ensure safety but significantly reduces helpfulness. Recent works have applied the RLHF for MLLMs to mitigate hallucinations in image understanding and mitigate trustworthiness. RLHF-V [37] introduced human preference through segment-level corrections for hallucination to enhance trustworthiness based on DDPO, a variant of DPO [27]. Similarly, Llava-RLHF [31] incorporated additional factual information into the reward model to reduce reward hacking and improve performance. However, these methods primarily focus on single-dimensional alignment for helpfulness without explicitly addressing safety concerns in MLLMs.
@@
Prior work has shown that efforts to enhance the helpfulness of MLLMs can inadvertently compromise their safety [1, 7]. Therefore, developing methods that improve helpfulness without undermining safety remains an open challenge. Inspired by Safe RLHF [13, 5], which decouples human preferences during data annotation. Through reward and cost models, it jointly optimizes the dual objective of helpfulness and safety. In this work, we propose Safe RLHF-V to enhance the helpfulness and safety of MLLMs. Our main contributions are as follows:
@@
  • (Dataset) We have open-sourced BeaverTails-V, the first dataset featuring dual preference annotations for both helpfulness and safety in MLLMs. For each pair, we independently annotated preferences regarding helpfulness and safety. Additionally, we provided graded safety labels: minor, moderate, and severe, for each Q-A pair, establishing the first exploration to enable multi-level safety alignment in MLLMs.
@@
  • (Algorithm) We introduce Safe RLHF-V, the first multimodal safety alignment algorithm that balances helpfulness and safety. This balance is achieved through a Lagrangian-based min-max optimization framework. Experimental results show that Safe RLHF-V significantly enhances model safety without sacrificing helpfulness. Specifically, Safe RLHF-V, the model’s safety improved by 34.2%, while helpfulness increased by 34.3%.
@@
Figure 1: Safety annotation distribution in the SPA-VL, UnsafeBench, and BeaverTails-V (ours) datasets. We quantified the distribution of images labeled as safe and unsafe (including minor, moderate, and severe) by expert annotations. This analysis reveals that existing datasets significantly overestimate the proportion of safe images. In contrast, BeaverTails-V presents a lower proportion of safe images, more accurately reflecting harmful categories.
@@ 2 Related Work
Alignment and Safety Concerns in MLLMs
Recent MLLMs [6, 36, 40, 35, 35, 15, 33] have raised safety concerns, including generate offensive content [17, 25], leak personal information [22, 10], and propagate misinformation [24]. Alignment aims to ensure these AI models are in accordance with human intention and values [11]. In practice, the widely accepted standard for large models is the 3H principle: Helpful, Honest, and Harmless [3]. Reinforcement Learning from Human Feedback (RLHF) [21] trains a reward model (RM) on a human preference dataset and uses the reward signal provided by the RM to fine-tune via reinforcement learning [30]. It is now widely applied to align LLMs and its variants [27]. Inspired by this, many methods that learn from human or AI preferences have been developed for aligning MLLMs [31, 37, 32]. LLaVA-RLHF [31] trains VLLMs and augments the reward model with additional factual information, mitigating reward hacking in RLHF. RLHF-V [37] further enhances MLLM trustworthiness through behavior alignment based on fine-grained corrective human feedback. Although these approaches have significantly improved the performance of MLLMs, they do not explicitly guarantee model safety, particularly when handling toxicity-related queries. To tackle this challenge, we introduce Safe RLHF-v, the first RLHF algorithm specifically designed to ensure the safety of MLLMs.
Safety Preference Data and Evaluation
Evaluating the safety of MLLMs has become a critical area of research [17, 23, 9, 34], especially as these models are increasingly deployed in real-world applications. Various benchmarks have been developed to assess the robustness of MLLMs against potential vulnerabilities. SPA-VL [38] focuses on how models adapt to safety preferences, but its images mainly depict safe scenarios and its image-text pairs are not well decoupled, limiting the model’s ability to fully leverage multimodal interactions. Unsafebench [26] evaluates task executions in hazardous scenarios, but lacks sufficient diversity in dangerous situations. MM-SafetyBench [16] assesses models’ understanding of safety preferences, but is more focused on evaluation than providing comprehensive training data. Although these datasets provide valuable information on multimodal safety preferences, they face limitations in data diversity, modality decoupling, and the balance between training and evaluation. To address these shortcomings, we propose Beavertails-V, aiming to offer a more diverse and comprehensive dataset for multimodal safety preferences.
@@ 3 Datasets
Datasets Models Blank Origin
SPA-VL LLaVA-1.5-7B 0.312 0.381 ($\uparrow$ 0.069)
Qwen2-VL-7B 0.029 0.107 ($\uparrow$ 0.078)
MM- LLaVA-1.5-7B 0.167 0.285 ($\uparrow$ 0.118)
Safetybench Qwen2-VL-7B 0.054 0.256 ($\uparrow$ 0.202)
Beavertails-V LLaVA-1.5-7B 0.455 0.554 ($\uparrow$ 0.099)
(Ours) Qwen2-VL-7B 0.083 0.412 ($\uparrow$ 0.329)
@@ 3.1 Why build a new safety MLLMs dataset?
The current open-source multimodal harmlessness datasets exhibit low consistency between text and image prompt inference. To investigate this, we conducted a correlation evaluation across BeaverTails-V, SPA-VL [38], and MM-SafetyBench [18], with results summarized in Tab. 1. For each harmful inference, two types of images were tested: (1) the original image relevant to the prompt and (2) a blank image. We utilized the general-purpose MLLM – LLaVA-1.5-7B [15] to perform adversarial attacks, measuring the attack success rate (ASR) [42, 18]. ASR is defined as the proportion of unsafe responses generated, and its calculation involves using the GPT-4O [1] API to assess the safety of the generated responses.
Figure 2: Example of harmful and unethical themes in BeaverTails-V: A collection of 20 panels categorized into 9 primary harmful domains, each labeled with a number (X-X), where the first X indicates the primary category and the second X represents the subcategory, covering a range of harmful and unethical behaviors. Warning: this paper contains example data that may be offensive or harmful.
Experimental results reveal that image content minimally influences the harmfulness of queries in existing datasets. For example, in SPA-VL, the ASR with a blank image was 31.2%, increasing marginally to 38.1% with the original image – an increment of only 6.9%. This outcome reinforces our hypothesis that current multimodal safety alignment datasets lack genuine multimodal integration.
Reasoning #2
Compared to existing multimodal safety datasets, BeaverTails-V offers unignorable advantages in the diversity of harm categories and well-refined annotations, as shown in Fig. 1. BeaverTails-V encompasses more harm categories with a balanced distribution between safe and unsafe data. For example, over 50% images in categories such as unfair behavior and high-risk financial activities are labeled as safe in SPA-VL. This imbalance limits the applicability of existing datasets in real-world scenarios.
Furthermore, BeaverTails-V achieves an orthogonal decoupling of helpfulness and safety while incorporating multi-level safety grading annotations to facilitate hierarchical safety alignment. Specifically, beyond categorizing QA pairs as safe or unsafe, BeaverTails-V further assigns meta-labels indicating the severity of harm (i.e. minor, moderate, and severe) to unsafe QA pairs. This detailed annotation framework fundamentally differentiates BeaverTails-V from existing open-source multimodal harmlessness datasets, enabling models to perform more precise and nuanced safety risk modeling.
@@ 3.2 Data Collection and Annotation Process:
To construct the high-quality BeaverTails-V dataset, we designed a multi-stage data collection and annotation process as described below.
@@
❶ Harmful Categories: Drawing inspiration from the text-based safety datasets [8, 13], and guided by usage policies and human assessment, we constructed a multimodal safety taxonomy with 9 primary and 20 secondary categories, as shown in Fig. 2. Detailed descriptions of each category are provided in Appendix.
❷ Image Collection: Leveraging our defined harm taxonomy, we used the Yandex search engine as the primary source for diverse image retrieval. GPT-4 generated category-aligned keywords, which were manually curated and refined to minimize discrepancies between retrieved images and category expectations. To balance representation across harm levels, we dynamically adjusted search strategies, intensifying retrieval for high-harm categories. All collected images underwent multiple rounds of manual filtering and validation to ensure quality and relevance.
Models Metrics Accuracy Precision Recall F1-Score↑ False Positive Rate↓
Llama-Guard-3-11B-Vision Safety 0.64 0.67 0.84 0.74 0.7077
Beaver-Guard-V (Ours) Safety 0.86 0.88 0.90 0.89 0.2131
Table 2: Comparison between Beaver-Guard-V and other methods.
Models $\downarrow$ FoN $\rightarrow$ @0 $\downarrow$ @1 $\downarrow$ @3 $\downarrow$ @5 $\downarrow$
LLaVA-1.5-7B 0.502 0.398 0.316 0.281
Qwen2-VL-7B 0.326 0.275 0.258 0.249
LLaVA-v1.6-13B 0.343 0.220 0.173 0.154
Table 3: Attack success rates under different moderation levels. The comparison of attack success rates for different VLMs as the moderation level increases.
❸ Prompt Generation: Generating prompts that are both meaningful and contextually relevant to the images is crucial for capturing the complexity of harmful content. To address this, we combined instruction-following and jailbreak strategies. Using the Ovis1.6-Gemma2-9B model, we generated eight prompt variations per image, striking a balance between diversity and quality.
Initial prompts were filtered to remove invalid content. From the remaining candidates, the most diverse and representative prompts were selected. An additional manual validation step further enhanced diversity. The final set of prompts underwent rigorous human review to ensure alignment with dataset objectives and semantic richness.
@@
Response Generation: We utilized generic VLLMs to generate responses for 32k prompts. To maximize diversity and representativeness in response pairs, we applied a two-step filtering strategy: (1) one response was randomly selected from five model outputs (excluding Phi-3.5-vision-instruct [19]) as response a. (2) using paraphrase-MiniLM-L6-v2 [28] embeddings and cosine similarity, we selected the most dissimilar response as response b.
Existing models often struggled to generate concise rejection responses, leading to out-of-distribution issues. To mitigate this, we incorporated rejection responses generated by the Phi-3.5-vision-instruct model at a 4% ratio. This embedding-based selection method effectively enhanced dataset diversity.
@@
❺ Question-Answer Pair Annotation: Prior works have shown that GPT-4o matches human annotators’ accuracy and preference consistency for data annotation tasks [39]. Inspired from PKU-SafeRLHF [12], we identified distinct gradations in the helpfulness and harmlessness of model-generated responses. While models can generate basic rejections, these are generally less safe than responses offering proactive warnings and detailed explanations.
We defined safety on a 7-point scale, with a score of 0 representing a neutral balance between safety and harm. Negative scores (-1 to -3) denote progressively increasing levels of harmfulness, whereas non-negative scores indicate ascending degrees of safety. In particular, scores of 0 and 1 correspond to basic rejections without guidance, while scores of 2 and 3 reflect progressively stronger advisory tones and responsibilities.

### v2

# Safe RLHF-V: Safe Reinforcement Learning from Multi-modal Human Feedback

@@ Abstract
Multimodal large language models (MLLMs) are essential for building general-purpose AI assistants; however, they pose increasing safety risks. How can we ensure safety alignment of MLLMs to prevent undesired behaviors? Going further, it is critical to explore how to fine-tune MLLMs to preserve capabilities while meeting safety constraints. Fundamentally, this challenge can be formulated as a min-max optimization problem. However, existing datasets have not yet disentangled single preference signals into explicit safety constraints, hindering systematic investigation in this direction. Moreover, it remains an open question whether such constraints can be effectively incorporated into the optimization process for multi-modal models. In this work, we present the first exploration of the Safe RLHF-V – the first multimodal safety alignment framework. The framework consists of: (I) BeaverTails-V, the first open-source dataset featuring dual preference annotations for helpfulness and safety, supplemented with multi-level safety labels (minor, moderate, severe); (II) Beaver-Guard-V, a multi-level guardrail system to proactively defend against unsafe queries and adversarial attacks. Applying the guard model over five rounds of filtering and regeneration significantly enhances the precursor model’s overall safety by an average of 40.9%. (III) Based on dual preference, we initiate the first exploration of multi-modal safety alignment within a constrained optimization. Experimental results demonstrate that Safe RLHF effectively improves both model helpfulness and safety. Specifically, Safe RLHF-V enhances model safety by 34.2% and helpfulness by 34.3%. [^1] All of datasets, models, and code can be found at https://github.com/SafeRLHF-V to support the safety development of MLLMs and reduce potential societal risks.
@@ 1 Introduction
Multimodal large language models (MLLMs) are pivotal for developing general-purpose AI assistants [1, 2]. By using visual instruction tuning on foundation models [3], MLLMs can effectively handle complex multimodal tasks. However, recent research highlights that images can implicitly induce MLLMs to generate harmful content, exposing vulnerabilities that are less pronounced in purely LLMs. What is worse, fine-tuning MLLMs can lead to the forgetting of safety alignment previously learned by the backbone LLMs [4]. Consequently, improving the safety of MLLMs to align with human preference is a major concern at present [5]. So, as MLLMs continue to advance, emerging studies reveal a critical challenge:
How can we ensure the safety alignment of MLLMs? This is under exploration.
So, we need to answer: Which type of preference data and what optimization are effective?
@@
Reinforcement Learning from Human Feedback (RLHF) has proven effective in aligning LLMs with human preferences [6, 7, 8, 9, 10], enhancing capabilities in instruction following, reasoning, etc. However, balancing helpfulness and safety remains challenging. Specifically, during the preference annotation stage, it is challenging to effectively account for two conflicting metrics within a single partial order. Helpfulness and safety are inherently somewhat contradictory; for instance, simply refusing to answer may ensure safety but significantly reduce helpfulness. Recent works have applied the RLHF for MLLMs to mitigate hallucinations in image understanding and mitigate trustworthiness. RLHF-V [11] introduced human preference through segment-level corrections for hallucination to enhance trustworthiness based on DDPO, a variant of DPO [7]. Similarly, Llava-RLHF [12] incorporated additional factual information into the reward model to reduce reward hacking and improve performance. However, these methods primarily focus on single-dimensional alignment for helpfulness without explicitly addressing safety concerns in MLLMs.
@@
Prior studies have shown that attempts to enhance the helpfulness of MLLMs can inadvertently compromise their safety [13, 14], without guaranteeing safety. Consequently, developing methods that enhance helpfulness without compromising safety remains an open challenge. Inspired by Safe RLHF [15, 16], which decouples human preferences during data annotation and jointly optimizes the dual objectives of helpfulness and safety. However, directly applying Safe RLHF to multimodal learning proves insufficient in practice. We identify several key challenges, as detailed below:
  • Lack of high-quality multimodal safety data. Existing multimodal safety preference datasets often exhibit weak correlations between visual and textual modalities. In particular, the presence of harmful or benign images tends to exert minimal influence on the prompt’s harmfulness. Overall, the images and text in current datasets appear largely independent, which limits their utility in learning grounded safety preferences.
  • Direct safety meta-label annotation is infeasible. In language-only settings, explicitly categorizing safety-related preferences has proven highly effective. However, the introduction of multimodal inputs introduces substantial inconsistencies in labeling. Annotators tend to focus on different aspects of an image, resulting in divergent interpretations and making it challenging to generate consistent and reliable safety preference labels.
  • Whether constrained optimization remains effective in the multi-modal setting remains to be verified. Although constrained optimization has been well-studied for language models, its effectiveness in the multi-modal context remains uncertain and poses significant engineering challenges. In our exploration, we found the training process in multi-modal highly unstable, with significant difficulty in diagnosing the root cause of failures. When model performance stagnates, it is often unclear whether the issue arises from flawed preference data, suboptimal training hyperparameters, or - perhaps most critically - the inherent ineffectiveness of constrained optimization in multi-modal settings.
In this work, we propose Safe RLHF-V, a novel framework designed to safety alignment of MLLMs, addressing the aforementioned challenges through innovations in data construction, annotation strategy, and optimization methodology. Our main contributions are as follows:
@@
  • (Dataset) We have open-sourced BeaverTails-V, the first dataset featuring dual preference annotations for both helpfulness and safety in MLLMs. For each pair, we independently annotated preferences regarding helpfulness and safety. Additionally, we provided graded safety labels: minor, moderate, and severe to reduce the inconsistencies in labeling, establishing the first exploration to enable multi-level safety alignment in MLLMs.
@@
  • (Algorithm) We introduce Safe RLHF-V, the first multimodal safety alignment algorithm that balances helpfulness and safety. This balance is achieved through a Lagrangian-based min-max optimization framework. To reduce the unstable optimization, we introduce the budget bound. Experimental results show that Safe RLHF-V significantly enhances model safety without sacrificing helpfulness. Specifically, Safe RLHF-V, the model’s safety improved by 34.2%, while helpfulness increased by 34.3%.
@@ 2 Related Work
Alignment and Safety Concerns in MLLMs.  Recent LLM [17, 18, 19, 20, 21] and MLLMs [22, 23, 2, 24, 24, 25] have raised safety concerns, including generate offensive content [21, 26], leak personal information [27, 28], and propagate misinformation [29]. Alignment aims to ensure these AI models are following human intention and values [5]. In practice, the widely accepted standard for large models is the 3H principle: Helpful, Honest, and Harmless [30]. Reinforcement Learning from Human Feedback (RLHF) [6] trains a reward model (RM) on a human preference dataset and uses the reward signal provided by the RM to fine-tune via reinforcement learning [31]. It is now widely applied to align LLMs and its variants [7]. Inspired by this, many methods that learn from human or AI preferences have been developed for aligning MLLMs [12, 11, 32]. LLaVA-RLHF [12] trains VLLMs and augments the reward model with additional factual information, mitigating reward hacking in RLHF. RLHF-V [11] further enhances MLLM trustworthiness through behavior alignment based on fine-grained corrective human feedback. Although these approaches have significantly improved the performance of MLLMs, they do not explicitly guarantee model safety, particularly when handling toxicity-related queries.
Safety Preference Data and Evaluation.  Evaluating the safety of MLLMs has become a critical area of research [21, 33, 34, 35], especially as these models are increasingly deployed in real-world applications. Various benchmarks have been developed to assess the robustness of MLLMs against potential vulnerabilities. SPA-VL [36] focuses on how models adapt to safety preferences, but its images mainly depict safe scenarios, and its image-text pairs are not well decoupled, limiting the model’s ability to fully leverage multimodal interactions. Unsafebench [37] evaluates task executions in hazardous scenarios, but lacks sufficient diversity in dangerous situations. MM-SafetyBench [38] assesses models’ understanding of safety preferences, but is more focused on evaluation than providing comprehensive training data. Although these datasets provide valuable information on multimodal safety preferences, they face limitations in balancing helpfulness and safety. To address these shortcomings, we propose Beavertails-V, aiming to offer a more diverse and comprehensive dataset for multimodal safety preferences.
@@ 3 Datasets
[…]
SPA-VL LLaVA-1.5-7B 0.213 0.401 0.188
Qwen2-VL-7B 0.040 0.108 0.068
MM- LLaVA-1.5-7B 0.041 0.233 0.192
Safetybench Qwen2-VL-7B 0.002 0.121 0.119
Beavertails-V LLaVA-1.5-7B 0.349 0.565 0.216
(Ours) Qwen2-VL-7B 0.036 0.276 0.240
@@ 3.1 Why build a new safety MLLMs dataset?
The current open-source multimodal harmlessness datasets exhibit low consistency between text and image prompt inference. To investigate this, we conducted a correlation evaluation across BeaverTails-V, SPA-VL [36], and MM-SafetyBench [39], with results summarized in Table 1. For each harmful inference, two types of input were tested: (1) the original text-image pair and (2) the text-only input. We utilized the general-purpose MLLM – LLaVA-1.5-7B [3] and Qwen2-VL-7B [25] to perform adversarial attacks, measuring the attack success rate (ASR) [40, 39]. ASR is defined as the proportion of unsafe responses generated, and its calculation involves using the GPT-4o [13] API to assess the safety of the generated responses. Experimental results reveal that image content minimally influences the harmfulness of queries in existing datasets. For example, in SPA-VL evaluated with Qwen2-VL-7B, the ASR with text-only input was 4.0%, increasing marginally to 10.8% with the original text-image pair – an increment of only 6.8%. This outcome reinforces our hypothesis that current multimodal safety alignment datasets lack genuine multimodal integration.
Reasoning #2 . Compared to existing datasets, BeaverTails-V offers unignorable advantages in the diversity of harm categories and well-refined annotations, as shown in Fig. 1. BeaverTails-V encompasses more harm categories with a balanced distribution between safe and unsafe data. For example, over 50% images in categories such as unfair behavior and high-risk financial activities are labeled as safe in SPA-VL. This imbalance limits the applicability in real-world scenarios. Furthermore, BeaverTails-V achieves an orthogonal decoupling of helpfulness and safety while incorporating multi-level safety grading annotations to facilitate hierarchical safety alignment. Specifically, beyond categorizing QA pairs as safe or unsafe, we further assign meta-labels indicating the severity of harm (i.e. minor, moderate, and severe) to unsafe QA pairs.
Figure 1: Safety annotation distribution of various multimodal safety dataset. We labeled the type of images as safe and unsafe (including minor, moderate, and severe) by expert annotations. We find that existing datasets significantly overestimate the proportion of safe images. In contrast, BeaverTails-V presents a lower proportion of safe images, more accurately reflecting harmful categories.
@@ 3.2 Data Collection and Annotation Process:
To construct the high-quality BeaverTails-V dataset, we designed a multi-stage data collection and annotation process as described below. A detailed pipeline document can be referred to Appendix B.
  • $\hookrightarrow$
@@
    Harmful Categories. Drawing inspiration from the text-based safety datasets [41, 15], and guided by usage policies and human assessment, we constructed a multimodal safety taxonomy with 9 primary and 20 secondary categories, as shown in Fig. 5. Detailed descriptions of each category are provided in Appendix B.2.
  • $\hookrightarrow$
    Prompt and Image Generation: We constructed our dataset by first retrieving diverse, harm-aligned images from Yandex, guided by a taxonomy-driven set of GPT-4 generated keywords that were manually refined to ensure category fidelity. To balance harm-level representation, we dynamically adapted search strategies and performed multi-round manual filtering for quality control. For each image, we then used the Ovis1.6-Gemma2-9B model to generate eight diverse and contextually rich prompts via a combination of instruction-following and jailbreak strategies, aiming to capture the nuanced semantics of harmful content.
  • $\hookrightarrow$
@@
    Response Generation: We utilized various VLLMs to generate responses for 32k prompts. To maximize diversity in response pairs, we applied a two-step filtering strategy: (1) one response was randomly selected from five model outputs (excluding Phi-3.5-vision-instruct [42]) as response a. (2) using paraphrase-MiniLM-L6-v2 [43] embeddings and cosine similarity, we selected the most dissimilar response as response b. Existing models often struggled to generate concise rejection responses, leading to out-of-distribution issues. To mitigate this, we incorporated rejection responses generated by the Phi-3.5-vision-instruct model at a 4% ratio. This embedding-based selection method effectively enhanced dataset diversity.
  • $\hookrightarrow$
@@
    Question-Answer Pair Annotation: Prior works have shown that GPT-4o matches human annotators’ accuracy and preference consistency for data annotation [44]. Inspired by PKU-SafeRLHF [45], we identified distinct gradations in the helpfulness and safety of model responses. While models can generate basic rejections, these are generally less safe than responses offering proactive warnings and detailed explanations. We defined safety on a 7-point scale, with a score of 0 representing a neutral balance between safety and harm. Negative scores (-1 to -3) denote increasing levels of harmfulness, whereas non-negative scores indicate ascending degrees of safety. In particular, scores of 0 and 1 correspond to basic rejections without guidance, while scores of 2 and 3 reflect stronger advisory tones and responsibilities.
Figure 2: Illustration of Beaver-Guard-V’s Moderation Pipeline. (a)&(b): Traditional moderation filters toxic prompts pre-generation, often refusing to respond, which can yield unhelpful or harmful outputs. Beaver-Guard-V adopts a two-stage approach, moderating both prompts and responses post-generation via rejection and regeneration to ensure safe, helpful outputs. (c): It provides fine-grained QA-pair annotations (Safe, Minor, Moderate, Severe) for downstream safety analysis. (d): Across multiple moderation rounds, Beaver-Guard-V significantly lowers attack success rates on models.
Table 2: Comparison of Beaver-Guard-V with baselines
2a Comparison between Beaver-Guard-V and other methods.
[…]
Llama-Guard-3-11B-Vision Safety 0.64 0.67 0.84 0.74 0.71
Beaver-Guard-V-Binary (Ours) Safety 0.78 0.77 0.93 0.84 0.47
Beaver-Guard-V-Multi-Level (Ours) Safety 0.85 0.88 0.87 0.88 0.20
2b Comparison between Beaver-Guard-V and Llama-Guard-3-11B-Vision.
BeaverTails-V SPA-VL MM-SafetyBench
Models Metrics Eval Test SD SD_TYPO TYPO Average
Llama-Guard-3-11B-Vision Accuracy 0.64 0.64 0.70 0.65 0.68 0.68
Beaver-Guard-V (Ours) Accuracy 0.85 0.88 0.88 0.85 0.90 0.88
Llama-Guard-3-11B-Vision F1-Score 0.74 0.74 0.80 0.74 0.79 0.78
Beaver-Guard-V (Ours) F1-Score 0.88 0.92 0.93 0.90 0.94 0.92